# Object as a Service: Simplifying Cloud-Native Development through Serverless Object Abstraction

### v1

@@ Abstract
Index Terms:
FaaS, Serverless paradigm, Cloud computing, Cloud-native programming, Abstraction.
@@ I-A FaaS and Its Shortcomings
Function-as-a-Service (FaaS) paradigm is becoming widespread and envisioned as the next generation of cloud computing systems (a.k.a. Cloud 2.0) [18] that mitigates the burden for programmers and cloud solution architects. Major public cloud providers offer FaaS services (e.g., AWS Lambda, Google Cloud Function, Azure Function), and several open-source platforms for on-premise FaaS deployments are emerging (e.g., OpenFaaS, Knative). FaaS offers the function abstraction that allows users to develop their business logic and invoke it via a predefined trigger. In the backend, the serverless platform hides the complexity of resource management and deploys the function seamlessly in a scalable manner. FaaS is proven to reduce development and operation costs via implementing scale-to-zero and charging the user in a truly pay-as-you-go manner. Thus, it aligns with modern software development paradigms, such as CI/CD and DevOps [7].
As the FaaS paradigm is primarily centered around the notion of stateless functions, it naturally does not deal with the data. However, in practice, most use cases must maintain some form of (structured or unstructured) state and keep them in the external data store. Thus, often the developers have to intervene and undergo the burden of managing the application data using separate cloud services (e.g., AWS S3 [4]). Even though stateless functions make the system scalable and manageable, the state still exists in the external data store, and the developer must intervene to connect the function to the data store. For instance, in a video streaming application [12], developers must maintain video files, metadata, and access control in addition to developing functions.
@@
Last but not least, current FaaS abstractions do not natively support function workflows. To form a workflow, the developer has to generate an event that triggers another function in each function. However, configuring and managing the chain of events for large workflows becomes cumbersome. Although function orchestrator services (e.g., AWS Step Function [3] and Azure Durable Function [9]) can be employed to mitigate this burden, the lack of built-in workflow semantics (see Figure 1) in FaaS forces the developer to intervene and employ other cloud services to navigate the data throughout the workflow manually. In sum, although FaaS makes the resource management details (e.g., auto-scaling) transparent from the developer’s perspective, it does not do so for the data, access control, and workflow management.
@@ I-B Proposed Paradigm
As shown in Figure 1, unlike FaaS, OaaS segregates the state management from the developer’s source code and incorporates it into the serverless platform to make it transparent from the developer’s perspective. OaaS also incorporates workflow orchestration via its dataflow abstraction with built-in data navigation across functions. Incorporating the application data and workflow into the object abstraction unlocks opportunities for built-in optimization features, such as data locality, data reliability, caching, software reusability [13], and data access control. OaaS provides an interface for developers to declare the behavior and properties of objects in the form of class and function. For more complex use cases where the developer needs to orchestrate multiple objects and functions in the form of a workflow, OaaS offers the native (built-in) dataflow semantic [37] that can hide details of the data navigation and synchronization via declaring the flow of data in the form of an object’s function.
An exemplar use case that can take advantage of OaaS is a cloud-based video streaming system (e.g., [12, 31]) that needs developers to implement new streaming services for the available video content rapidly. Examples of such services are blurring harmful (gore) content from videos, detecting faces on surveillance videos, and making the video processing pipeline for transcoding. Implementing these services using FaaS entails dealing with the state data (i.e., videos), in addition to developing the business logic. In this scenario, the OaaS paradigm can mitigate the developer’s burden by offering encapsulation. The videos are defined as persistent objects bound to a set of functions that can be invoked by the viewer’s application and potentially change the object’s (video’s) state. For instance, to blur gore segments in the video object v1, the developer invokes v1.detect_gore() to get output object g1 that contains a list of time stamps of the gore segments. Next, blurring the video is performed via invoking v1.blurs(g1).
@@ I-C Research and Contributions
The first and foremost challenge to establish OaaS is to offer object abstraction with an interface for developers to declare the behavior and properties of objects in the form of class, function, and dataflow semantics. Together with access modifiers, OaaS can hide the internal mechanics (i.e., private function) from other developers and users. The object can have pointers to other objects and form a dataflow function on top of them. Such a high-level object can hide the details of low-level ones by declaring access modifiers that reject any function calls that involve invalid access.
To realize the OaaS paradigm, we develop the Oparaca (Object Paradigm on Serverless Cloud Abstraction) platform with the concern of the flexibility and modularity, so that it can accommodate different use cases. The second challenge is to design Oparaca to be pluggable and can work with different execution (i.e., FaaS) and storage modules. We avoid having Oparaca tightly depend on the execution module via exploiting the notion of pure function that offloads the task to the execution module without any side-effect. We also devise a data tiering scheme as the abstraction layer between the storage and execution modules. Therefore, changing the storage type does not imply changing the function code.
@@
Although utilizing the aforementioned invocation scheme provides Oparaca with ease of use and flexibility, it causes additional overheads, mainly because of moving the data between OaaS modules. This is particularly important for unstructured data, also known as BLOB (e.g., multimedia content), that usually persists on a separate storage type, such as object storage. Therefore, the third challenge is to design Oparaca so that supporting objects’ overhead is minimal and tractable. To mitigate the overhead, the data tiering scheme within the Oparaca platform diminishes the latency of accessing the object by using distributed in-memory hash table [40]. To further mitigate the overhead of accessing objects, Oparaca is equipped with a presigned URL with a redirection mechanism that reduces unnecessary data movements within the platform instead of relaying (transferring) the object state.
@@
In OaaS, concurrent access requests can be made to the same state, which makes it prone to data inconsistency from race conditions. Moreover, any system failure can potentially cause state inconsistency between multiple data stores. For example, if the function updates the state to the first storage but fails to update the second storage, it may cause an inconsistency. Thus, the fourth challenge is to make Oparaca fault-tolerant and maintain data consistency across multiple data stores. To overcome this challenge, we develop a fail-safe state transition in Oparaca to maintain consistency across various data stores. In addition, we develop a method based on the lightweight optimistic locking [29] together with localized locking to avoid concurrent same-object modification. Another aspect of fault tolerance is the recovery mechanism from failure via retrying. However, such a remedy by itself can lead to another problem—repeating the execution more than once and falling into an undesirable state. To cope with this challenge, we develop a mechanism within Oparaca that establishes consistent state transitioning by guaranteeing “exactly-once” execution for the function calls.
@@
    Proposing the OaaS paradigm to hide data, workflow, and resource management complexities from the user view, and to mitigate the redundancy complexity through developing the cloud object inheritance mechanism.
@@
    Developing a prototype of OaaS, called Oparaca platform [^1] The source code, packages, and example use cases are available here: https://github.com/hpcclab/OaaS, that can support both structured and unstructured types of states while maintaining data consistency and fault-tolerance with minimal performance overhead.
@@
    Minimizing the overhead of Oparaca via presigned URL with redirection and caching mechanisms.
@@
In the rest of this paper, Section II reviews the state of the art in the serverless paradigm. Section III discusses the conceptual design. Section IV presents the architecture of Oparaca and internal details. Section V discusses non-focused topics in this work. Section VI evaluates the overhead, scalability, and development efficiency of OaaS. Finally, we conclude this paper in Section VII.
@@ II Background and Prior Studies
II-A FaaS Paradigm
The FaaS paradigm allows the developer to implement the application as a set of independent functions that are transparently provisioned in isolation on the cloud infrastructure. FaaS is offered by public cloud providers (e.g., AWS Lambda, Azure Function, Google Cloud Function).
FaaS can also be self-hosted via open-source platforms (e.g., OpenFaaS [14] and OpenWhisk [15]). FaaS invokes the function that matches its predefined trigger(s) upon receiving the event.
A variant of FaaS, Container as a Service (CaaS) [20], does not offer the function development framework. Instead, the user must provide the already-containerized function. Kubernetes [10] is a widely-used platform that automates container provisioning (via YAML configuration language) and manages the life-cycle of containerized services. Knative [16] complements Kubernetes by enabling FaaS features at the container level. It is composed of two main components: Knative Serving, and Knative Eventing. The former enables auto-scaling, scale to zero, and minimal configuration of the containerized services. The latter enables pipelining and routing events to streamline the development of event-driven containers.
II-B State Management Strategies
The idea of stateful serverless is explored in several research works. As noted in Figure 2 and Table I, these works can be categorized into actor model, datastore abstraction, and pure function approaches depending on where the platform stores the state data, and how the function accesses the data. According to Figure 2, the actor model places the state inside the worker node (instance) to achieve the data locality. In the pure function, however, the state is placed on other systems (e.g., database) and is transferred to the worker instance upon invocation. Hence, the state appears as part of the function input argument, and the modified state appears as its output. Thus, the function is still stateless while exhibiting stateful features. The datastore abstraction is a hybrid approach where the platform provides the API for the function to access the data on demand. Depending on the design, the state might be stored in the database but can also be cached in the worker.
@@ II-B State Management Strategies
Actor
Model
Datastore
Abstraction
Pure
Function
Data
placement
worker
instances
platform services
or database
(depend on impl.)
platform
services
or database
Complexity high depend on impl. low
Data locality high depend on impl. low
Unstructured
data support
difficult yes difficult
Deployment
granularity
actor
(multiple
functions)
function function
Maintain-
ability
low high high
Solutions
Kalix [22],
Azure Entity
Func.[33],
Cepless [32]
Cloudburst [38],
FAASM [35],
Apiary [27],
Crucial [8],
Boki [25],
Beldi [44],
Oparaca
Kalix [22]
Statefun [5],
Oparaca
TABLE I: Comparing properties of various design patterns to build a stateful serverless platform.
According to Table I, the actor model serverless platform needs to maintain the availability of each actor (object) where both data and compute reside. Maintainability is particularly difficult for bulky unstructured data because the platform needs to balance each node’s computing and storage aspects. In addition, the platform has to support a routing mechanism to navigate a function call to the actor’s location. Alternatively, the pure function approach disaggregates the state management and compute (function) for the sake of system design simplicity. However, it compromises the data locality aspect. Similar to pure functions, datastore abstraction also relaxes the need to store the state on the worker node. Regardless, it utilizes caching techniques to preserve data locality. The deployment granularity of the actor model approach is an actor with multiple functions that share the same state, whereas the granularity in other approaches is a single function.
The actor model approach has been popular in programming languages and OOP, because it spurs asynchronous messaging across actors and lends itself to distributed deployments. That is why it has been an attractive choice for stateful serverless platforms, even though it poorly supports unstructured data. Kalix [22] and Azure Entity Functions [33], which are part of Azure durable functions, are example platforms implemented based on the actor model approach. The serverless platforms based on the datastore abstraction are mostly popular in research. Cloudburst [38] offers stateful functions using a shared distributed key-value database to keep track of the state. FAASM [35] optimizes the function-state interaction overhead via employing web assembly [17]. Even though web assembly enables multiple functions to share the memory and achieve data locality, it implies compiling the code into web assembly, which limits the compatibility of some libraries. Crucial [8] allows a function to access the shared data via the DSO layer (distributed hash table). Boki [25] enables stateful function via providing API access to the distributed logging system. Beldi [44], on the other hand, provides the database and transaction API to the state. Apache Flink Stateful Function (StateFun) [5], and Kalix [22] are solutions based on the pure function approach. Statefun is built atop Apache Flink based on the actor model. However, it offloads the function code to a dedicated worker that can be categorized under the pure function approach.
@@ III-A Conceptual Modeling of OaaS
Third, OaaS offers built-in access control to provide the ability to declare the “scope of accessibility” for a state or function. Importantly, to define a set of related classes, the developer can declare it within a single module (a.k.a. package) that includes the access modifier to prevent unauthorized access to the state(s)/function(s) from other packages. This is particularly useful when cloud application developers utilize imported third-party packages.
Fourth, to enable higher-level abstractions for the users or developers, OaaS allows combining (nesting) cloud objects into one. The high-level object holds a reference to the lower-level object(s), and the invoked function can leverage the reference to fetch the lower-level object as its input. Importantly, we make it possible that the high-level object implements a dataflow function (called macro function); i.e., , invoking a chain of functions from the lower-level objects. This resembles configuring a workflow in conventional FaaS [3], albeit in a transparent manner. The major difference between macro functions and other FaaS workflows is that macros introduce the flow of execution via the flow of data rather than the invocation order. That is, given the dataflow semantics, the execution flow in a macro function is determined by the flow of data, and the developer only needs to introduce the flow. Then, in the background, OaaS handles the parallelism and data navigation while guaranteeing state consistency.
@@
III-B Developing Classes in OaaS
Figure 3: Different types of functions supported by OaaS.
@@ IV-A Design Goals
Now that the notion of OaaS is established, we elaborate on the design goals and architecture of the Oparaca platform. We set two design goals for Oparaca:
The first goal considers the fact that even though OaaS provides a higher-level abstraction for cloud-native application development, it is not always a replacement for FaaS. Hence, Oparaca should be backward compatible so that the developer can still use stateless FaaS and retrieve the data from the storage system directly. Additionally, most FaaS solutions support both synchronous and asynchronous function invocations. Therefore, Oparaca should support both types of function invocations as well for backward compatibility.
Second, for the sake of extensibility, Oparaca separates the components of the control plane from the execution plane, and the execution plane can be performed without the knowledge of its control plane but communicate through the standardized APIs. In this manner, the design is platform-agnostic and is not bound to a single preset underlying FaaS system, as its execution plane. This enables Oparaca to accommodate various types of execution planes, optimized for the requirements of different use cases, such as supporting latency-constrained function calls [36]; or supporting access to hardware accelerators [41].

### v2

@@ I-A FaaS and Its Shortcomings
Function-as-a-Service (FaaS) paradigm is becoming widespread and envisioned as the next generation of cloud computing systems (a.k.a. Cloud 2.0) [19] that mitigates the burden for programmers and cloud solution architects. Major public cloud providers offer FaaS services (e.g., AWS Lambda, Google Cloud Function, Azure Function), and several open-source platforms for on-premise FaaS deployments are emerging (e.g., OpenFaaS, Knative). FaaS offers the function abstraction that allows users to develop their business logic and invoke it via a predefined trigger. In the backend, the serverless platform hides the complexity of resource management and deploys the function seamlessly in a scalable manner. FaaS is proven to reduce development and operation costs via implementing scale-to-zero and charging the user in a pay-as-you-go manner. Thus, it aligns with modern software development paradigms, such as CI/CD and DevOps [9].
As the FaaS paradigm is primarily centered on stateless functions, it naturally does not deal with data. However, in practice, most use cases must maintain some form of state data. Moreover, FaaS limits direct communication between functions, making the application require additional remote storage for intermediate data. Thus, developers often have to intervene and undergo the burden of managing data using separate cloud services (e.g., AWS S3 [7]). For instance, in a video streaming application [14], developers must maintain video files, metadata, and access control in addition to developing functions.
@@
Last but not least, current FaaS abstractions do not natively support function workflows. To form a workflow, the developer has to generate an event that triggers another function in each function. However, configuring and managing the chain of events for large workflows is cumbersome. Function orchestrator services (e.g., AWS Step Function[6] and Azure Durable Function [11]) can be employed to mitigate this burden. However, FaaS suffers from two inherent shortcomings: (a) limited direct communication between functions and (b) lack of built-in state management (see Figure 1). These shortcomings force developers to intervene and employ other storage services to keep the intermediate data and manually navigate the data throughout the workflow [31]. In sum, although FaaS makes the resource management details (e.g., auto-scaling) transparent from the developer’s perspective, it does not do so for data, access control, and workflow management.
@@ I-B Proposed Paradigm
As shown in Figure 1, unlike FaaS, OaaS segregates state management from the developer’s source code and incorporates it into the serverless platform, making it transparent from the developer’s perspective. OaaS also incorporates workflow orchestration via its dataflow abstraction with built-in data navigation across functions. Incorporating the application data and workflow into the object abstraction unlocks opportunities for built-in optimization features, such as data locality, data reliability, caching, as well as developer-guided hints to ensure Quality-of-Service targets [28], enhance software reusability [15], and data access control. OaaS provides an interface for developers to declare the behavior and properties of objects in the form of class and function. For more complex use cases where the developer needs to orchestrate multiple objects and functions in the form of a workflow, OaaS offers the native (built-in) dataflow interface that can hide details of the data navigation and synchronization via declaring the flow of data in the form of an object’s function.
An exemplar use case that can take advantage of OaaS is a cloud-based video streaming system (e.g., [14, 30]) that needs developers to rapidly implement new streaming services for the available video content. Examples of such services are blurring harmful (gore) content from videos, detecting faces on surveillance videos, and making the video processing pipeline for transcoding. Implementing these services using FaaS entails dealing with the state data (i.e., videos) and developing the business logic. In this scenario, the OaaS paradigm can mitigate the developer’s burden by offering encapsulation. The videos are defined as persistent objects bound to a set of functions that can be invoked by the viewer’s application and potentially change the object’s (video’s) state. For example, to blur gore segments in the video object v1, the developer invokes v1.detect_gore() to obtain the output object g1 that contains a list of time stamps of gore segments—next, blurring the video via invoking v1.blurs(g1).
@@ I-C Research and Contributions
The first challenge is to establish OaaS by providing the “object” abstraction while ensuring the platform’s modularity and extensibility. The primary requirement is to offer an interface for developers to declare object behavior and state as classes, functions, and dataflow. With the addition of access modifiers, OaaS can encapsulate internal mechanics, allowing objects to reference other objects and form dataflow functions. This high-level abstraction enforces strong encapsulation, preventing invalid access and thereby unraveling complexity. To realize the OaaS paradigm, we develop Oparaca (Object Paradigm on Serverless Cloud Abstraction), that is driven by flexibility and modularity. It must accommodate multiple use cases, remain extensible, and integrate seamlessly with various execution (FaaS) and storage modules. By leveraging pure functions, Oparaca avoids tight coupling with execution modules and offloads tasks without side effects. A data tiering scheme further abstracts storage, ensuring that changes in storage types do not necessitate modifications in the function code.
@@
The second challenge is to reduce the overhead of data movement between functions and the platform’s components. For that purpose, the data tiering scheme within the Oparaca platform diminishes the latency of accessing the object by employing a distributed in-memory hash table [37]. However, for unstructured data, also known as BLOB (e.g., multimedia content), that cannot fit in memory, Oparaca persisted them in a separate object storage. To further mitigate the overhead of accessing objects, Oparaca is equipped with a presigned URL with a redirection mechanism that reduces unnecessary data movements within the platform instead of relaying (i.e., transferring) the object state.
@@
The third challenge is to ensure fault tolerance and data consistency. Concurrent requests to the same state can lead to data inconsistencies due to race conditions, and system failures may result in mismatched states across multiple data stores. Oparaca introduces a fail-safe state transition mechanism that maintains data consistency and fault tolerance to tackle these issues. Additionally, Oparaca combines lightweight optimistic locking [27] and localized locking to prevent concurrent modifications to the same object. Another aspect of fault tolerance is the recovery mechanism from failure via retrying. However, such a remedy can potentially lead to another problem—repeating the execution more than once and falling into an undesirable state. To cope with this challenge, we develop a mechanism within Oparaca that establishes consistent state transitioning by guaranteeing “exactly-once” execution for the function calls.
@@
    Proposing A novel OaaS paradigm that simplifies cloud-native development by unifying application data, business logic, and workflows into a single object abstraction. Uniquely, OaaS applies principles from OOP to the serverless model by introducing cloud class inheritance, which enhances software reusability across various cloud functions and runtimes.
@@
    Developing Oparaca [^1] The source code, packages, and example use cases are available here: https://github.com/hpcclab/OaaS, as a prototype implementation of OaaS. Oparaca supports both structured and unstructured data while ensuring data consistency through a fail-safe state transition mechanism and a scalable, localized locking scheme to prevent race conditions without the overhead of cluster-wide coordination
@@
    Devising a declarative dataflow abstraction that allows developers to define complex function workflows based on data dependencies rather than task dependencies. The Oparaca platform makes this high-level abstraction practical by using efficient techniques, such as presigned URLs with redirection and distributed in-memory caching, to minimize performance overhead and improve scalability compared to traditional FaaS state management.
@@ II Background and Prior Studies
The idea of stateful serverless is explored in several research works. These approaches primarily aim to address the stateless nature of the FaaS model, where the burden of managing application data, access control, and function workflows is often shifted to the developer through separate cloud services. As noted in Figure 2, these works can be categorized into actor model, datastore abstraction, and pure function approaches depending on how the platform manages data and allows functions to access it.
@@
Actor Model. In the actor model, the platform fetches the state from persistent storage and places (i.e., caches) it inside a worker instance, then dispatches the request to where the state resides to achieve data locality. This approach, however, can make maintainability difficult for bulky unstructured data. The deployment granularity is an actor that contains multiple functions sharing the same state. The foundational platform in this space is Orleans [10], which introduces “virtual actors”. Its influence is evident in modern platforms such as Kalix [24] and Azure Entity Functions [17]. While effective for certain use cases, this model can tightly couple state and compute within a single language and environment. In contrast, Oparaca manages the object abstraction at the platform level, allowing functions implemented in different languages or runtimes to operate on an object, thereby offering greater flexibility.
Datastore Abstraction. The datastore abstraction is a hybrid approach where the platform provides an API for the function to access data on demand. Like pure functions, it relaxes the need for state and function co-location, but can utilize caching to preserve data locality. Several systems utilize this pattern. Cloudburst [36] uses a shared distributed key-value database, while Boki [23] provides API access to a distributed logging system. This log-based approach for consistency differs from Oparaca’s mechanisms, which use a combination of fail-safe state transitions for unstructured data and localized locking to handle race conditions. Beldi [41], on the other hand, provides the database and transaction API to the state. While providing direct transactional APIs offers strong guarantees, Oparaca abstracts these concerns from developers through its fail-safe versioning scheme and guarantees of exactly-once execution, aiming for a higher-level resilience model. FAASM [34] optimizes the function-state interaction by using WebAssembly [18]. Although this enables multiple functions to share memory and achieve data locality, it requires compiling code to WebAssembly, which can limit library compatibility. Oparaca’s hybrid model avoids this limitation by combining the pure function and datastore abstraction approaches, offering broader language support.
Pure Function. In the pure function approach, the state is placed on another system and transferred to the worker instance upon invocation, appearing as part of the function’s input arguments. This disaggregates state management and computation for system design simplicity, but can compromise data locality. Apache Flink Stateful Function (StateFun) [16] and Kalix [24] are solutions based on this approach. Oparaca combines the pure function approach for structured data with a lazy-fetch mechanism for unstructured data, practically employing both the pure function and datastore abstraction approaches.
Process Abstractions. While OaaS simplifies development by increasing abstraction, a competing trend in recent work involves stepping back from abstraction to grant developers more control for performance. This trend of using OS-inspired abstractions is evident in various contexts. For serverless platforms, Process-as-a-Service (PaaS) [12] introduces a cloud process to optimize execution, while for traditional always-on services, Nu [33] utilizes migratable logical processes.
Serverless Workflow and Dataflow Beyond state management for individual functions, composing them into complex applications is another significant challenge. Commercial orchestrators, such as AWS Step Functions and Azure Durable Functions, can mitigate the burden of chaining events; however, developers are often still required to navigate data between steps manually. The research community has proposed more integrated solutions. For instance, DataFlower [31] introduces a dataflow paradigm specifically for serverless workflow orchestration to optimize data movement between functions. Netherite [11] focuses on the efficient execution of these workflows through techniques like partitioned state management and collocated execution. While these systems significantly advance workflow performance and orchestration, OaaS approaches the problem from a different angle by integrating workflow directly into its core object abstraction.
Oparaca’s dataflow functions allow developers to declaratively define a workflow as a DAG based on data dependencies rather than task dependencies. The platform transparently manages parallelism, data navigation, and state consistency, abstracting these complexities from the developer.
@@ III-A Conceptual Modeling of OaaS
Third, OaaS offers built-in access control to provide the ability to declare the “scope of accessibility” for a state or function. Importantly, when defining a set of classes, the developer can declare it within a single package that includes the access modifier to prevent unauthorized access from other packages. This is particularly useful when cloud application developers utilize imported third-party packages.
Fourth, OaaS enables higher-level abstractions by allowing cloud objects to be nested, where a high-level object references lower-level objects. Functions can use these references to fetch inputs or invoke dataflow functions (called macro functions) that chain operations across lower-level objects. Unlike traditional FaaS workflows [6], macro functions determine execution flow based on dataflow rather than task (i.e., function call) dependency. Developers only define the data flow, while OaaS manages parallelism, data navigation, and state consistency transparently.
III-B Developing Classes in OaaS
Figure 3: Different types of functions supported by OaaS.
@@ IV-A Design Goals
The Oparaca platform is designed with its foundational goal of providing object abstraction with two additional design goals: backward compatibility and extensibility. first, while OaaS simplifies cloud-native application development, it is not always a replacement for FaaS; thus, Oparaca supports stateless FaaS and direct data access to storage systems. Second, for extensibility, Oparaca decouples the control plane from the execution plane, allowing the execution plane to operate independently via standardized APIs. This platform-agnostic design accommodates various execution planes optimized for specific use cases, such as latency-constrained function calls [35] or access to hardware accelerators [38].